OpenStudy (anonymous):

How do you find all the solutions for x in lx-8l=21

OpenStudy (anonymous):

x - 8 = 21 or -(x-8) = 21 solve these two equations

OpenStudy (anonymous):

lx-8l=21 ==> x-8 =21 ==> x=29 or x-8 =-21 ==> x=-13